Andrews Company has $95,000 available to pay dividends. It has 2,000 shares of 10%, $100 par, preferred stock and 20,000 shares of $10 par common stock outstanding. The preferred stock is selling for $125 per share, and the common stock is selling for $ 20 per share. Required Compute the dividend yield on the preferred stock and the common stock. If required, round your answers to one decimal place. Dividend yield Preferred stock % Common stock %
Answer #1

Dividend yield of preference:-

=(100*10%)/125

=8.00%

Dividend yield of common stock:-

Earnings available after paying preference =95000-(2000*10) =75000

Per share =75000/20000 =3.75

Dividend yield =3.75/20 =18.75%